OpenStudy (anonymous):

lim ((1/(x+h))-(1/x))/h h->0

OpenStudy (anonymous):

guess you are trying to find the derivative of \[f(x)=\frac{1}{x}\] right?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

i know the answer is 1/x^2 but we need to use that formula and solve it

OpenStudy (anonymous):

actually it is \[-\frac{1}{x^2}\]

OpenStudy (anonymous):

you just have to do the algebra \[\frac{1}{x+h}-\frac{1}{x}=\frac{x-(x+h)}{x(x+h)}=\frac{-h}{x(x+h)}\]

OpenStudy (anonymous):

now divide all this mess by h, so the h cancels in the numerator to give you \[\frac{-1}{x(x+h)}\] and finally take the limit as h goes to zero and you are done

OpenStudy (anonymous):

oh wow that was easy thank you

OpenStudy (anonymous):

yw